Can somebody help please?
Solve 28(132(39x - 20)) = 12481
I don't know where to start!

Answer by checkley77(12844)

You start inside the (( )) & work your way out.
28(132(39x - 20)) = 12,481
28(5,148x-2,640)=12,481
144,144x-73,920=12,481
144,144x=12,481+73,920
144,144x=86,401
x=86,401/144,144
x~.59940 ans.